New at this..clear steps would be great.

A compound of mercury and oxygen is heated in order to decompose the

compound. a 4.08g sample of mercury oxide upon heating gives off oxygen gas.

After heating the mass of mercury metal left behind was 3.78g.

calculate empirical formula of the compound.
Answer #1

1)Mass of oxygen = Total mass - Mass after heating

= 4.08 g - 3.78 g = 0.3 g oxygen.

2) Moles of oxygen = Mass (g) / Molar mass

= 0.3 g / 32 g/mol

= 0.009375 mole.

3) Moles of Mercury = 3.78 g / 200.6 g/mol

= 0.01884 mole.

4) Now, 0.009375 mole is small value and we divide by this value to both values as

Oxygen = 0.009375 / 0.009375 = 1

Mercury = 0.01884 / 0.009375 = 2.0

5) That is in empirical formula of mercury oxide contains

1 oxygen and 2 mercury.

Therefore, empirical formula is Hg2O.